A research group conducted an extensive survey of 2896 wage and salaried workers on issues ranging

1.

A research group conducted an extensive survey of 2896 wage and salaried workers on issues ranging from relationships with their bosses to household chores. The data were gathered through hour-long telephone interviews with a nationally representative sample. In response to the question, "What does success mean to you?" 1517 responded, "Personal satisfaction from doing a good job." Let p be the population proportion of all wage and salaried workers who would respond the same way to the stated question. Find a 90% confidence interval for p. (Round your answers to three decimal places.) 

lower limit = _______ 

upper limit = _______ 


2.

A research group conducted an extensive survey of 3019 wage and salaried workers on issues ranging from relationships with their bosses to household chores. The data were gathered through hour-long telephone interviews with a nationally representative sample. In response to the question, "What does success mean to you?" 1603 responded, "Personal satisfaction from doing a good job." Let p be the population proportion of all wage and salaried workers who would respond the same way to the stated question. How large a sample is needed it we wish to be 95% confident that the sample percentage of those equating success with personal satisfaction is within 3.0% of the population percentage? (Mint: Use P+ 0.53 as a preliminary estimate. Round your answer up to the nearest whole number.) 

_______ workers 
Answer #1

1)

sample success x = 1517
sample size n= 2896
pt estiamte p̂ =x/n= 0.5238
for 90 % CI value of z= 1.645 from excel:normsinv((1+0.9)/2)
margin of error E=z*std error   = 0.0153
lower bound=p̂ -E                       = 0.509
Upper bound=p̂ +E                     = 0.539

2)

here margin of error E = 0.030
for95% CI crtiical Z          = 1.960 from excel:normsinv((1+0.95)/2)
estimated prop.=p= 0.530
sample size n= p*(1-p)*(z/E)2= 1064 (rounding up)